Grade 12th passMechanics

In the bohr model, the electron of a hydrogen atom moves in a circular orbit of radius 5.3 × 10^-11m with a speed of 2.2 × 10^6ms^-1. Determine it's frequency and the current.

Leo
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
As we know
v = 2\pi r * f
now
use v = 2.2 * 10^6 m / sec
r = 5.3 * 10^-11 m
hence find f which is frequency in rev / sec
f will come out as 6.6 * 10^15 rev / sec